Solve the following equation for R1:
1/R= 1/R1 + 1/R2

Answer by checkley77(12844)

1/R= 1/R1 + 1/R2
1/R1=1/R-1/R2
1/R1=(R2-R)/R*R2
R1(R2-R)=R*R2
R1=R*R2/(R2-R) ANS.